OpenStudy (anonymous):

Below are the steps to solve an equation step 1: |x-2| +3=7 step 2: |x-2|=7-3 step 3: |x-2|=4 which of the following is correct step to solve the equation?

OpenStudy (anonymous):

x – 4 = –7

OpenStudy (study_together):

I think it's step 2.

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

x+2=4 -x-2=4 x+2=4 x-2=-4

OpenStudy (anonymous):

i need to find the correct step after

zepdrix (zepdrix):

For absolute value, we can drop the absolute bars by throwing a plus/minus symbol in front of the stuff.\[\large\rm |\color{royalblue}{stuff}|=4\]\[\large\rm \pm\color{royalblue}{stuff}=4\]Let's write the plus/minus on the other side though, that will make things easier.\[\large\rm \color{royalblue}{stuff}=\pm 4\]

zepdrix (zepdrix):

So notice that we actually have `two` equations now. One for the plus case, and another for the negative.\[\large\rm \color{royalblue}{stuff}=4\]\[\large\rm \color{royalblue}{stuff}= -4\]

OpenStudy (anonymous):

yeah is step 2

zepdrix (zepdrix):

And for our problem, the `stuff` is simply `x-2`.

zepdrix (zepdrix):

i guess everyone else got confused by the question lol

OpenStudy (anonymous):

ahh sorry

zepdrix (zepdrix):

Do you understand what I did with the absolute there? :O Your `stuff` is `x-2`. So you're looking for an option that matches one of these,\[\large\rm \color{royalblue}{x-2}=4\]\[\large\rm \color{royalblue}{x-2}=-4\]

OpenStudy (anonymous):

so its x-2=-4 would be the next step to solve the equation

zepdrix (zepdrix):

Option 4? Yay good job \c:/

OpenStudy (anonymous):

Thank you so much! algebra is kicking my butt this summer